OpenStudy (anonymous):

The equation of the line tangent to the curve y = (kx+8)/ (k+x) at x=-2 is y=x+4. What is the value of k?

OpenStudy (campbell_st):

differentiate using the quotient rule to find the equation of the gradient dy/dx = ((k + x)1 - ( k+x)k)/(k +x)^2 its known from the information the gradient is m = 1 when x = - 2 then 1 = (( k - 2) - (k - 2)k)/(k - 2)^2 simplifying 1 = -(k -2)(k - 1)/(k-2)^2 1 = (k - 1)/(k - 2) or k - 2 =- k + 1 2k = 3 k = 3/2 the value of k is 3/2

OpenStudy (anonymous):

Well this was a multiple choice question and i wanted to know the steps but 3/2 isnt an answer choice....the choices are -3, -1, 1, 3, and 4 and he answer is 3

OpenStudy (anonymous):

but im not sure whats wrong with ur work

OpenStudy (campbell_st):

its in the derivative it should be dy/dx = ((k + x)k) - (kx + 8)1)/(k + x)^2 dy/dx = (k^2 - 8)(k + x)^2 sorry about that

OpenStudy (anonymous):

oh no its ok.....can u kind of explain though why u found the derivative and what not

OpenStudy (campbell_st):

oops dy/dx = (k^2 - 8)(k +x)^2

OpenStudy (anonymous):

wait shud it be a divison sign btwn the 2 parenthesis....

OpenStudy (campbell_st):

the deriviative gives the equation of the gradient of the tangent. By substituting the x value, x = -2 you will get the gradient at a specific point on the curve derivative is dy/dx = (k^2 - 8)/(k + x)^2

OpenStudy (campbell_st):

then tangent equation has a gradient of 1 so evaluate the derivative dy/dx = 1 when x = -2 to find k so 1 = (k^2 - 8)/(k - 2)^2 or (k - 2)^2 = k^2 - 8 expanding gives then k^2 - 4k + 4 = k^2 - 8 -4k = -12 k = 3

OpenStudy (anonymous):

THANKS SO MUCH....I APPRECIATE IT =)

OpenStudy (anonymous):

especially u explaining it and showing it step by step....thanks 4 teh patience